# Hardware Trends Impacting Floating-Point Computations In Scientific Applications

### v1

@@ II Benchmarks
Throughout this paper, we will refer to several community benchmarks that have emerged over time, each serving a critical role in evaluating and exposing performance characteristics and limitations of the underlying hardware, as well as serving as a readily conveyed, widely understood record of progress. These benchmarks have become standard tools in the high-performance computing (HPC) community for assessing the efficiency and effectiveness of various computing systems. While in this paper, we focus on floating-point operation-focused benchmarks, other benchmarks exist. One such example is the Graph 500 [4], which measures a system’s performance on graph-based problems important for large dataset analysis.
@@
A more recent addition, HPL-MxP (formerly called HPL-AI) [10], or HPL for mixed precision, is designed to benchmark systems optimized for AI and machine learning workloads. By focusing on mixed-precision operations, HPL-MxP reflects the growing need for systems capable of efficiently handling lower-precision calculations typical in AI models, exposing the performance capabilities of modern hardware in these emerging domains without abandoning the needs of scientific computing.
@@ III-B The Co-Processor Era
While the 8087 might be the best-known example of this technology, co-processors were in widespread use at the time. For example, the Motorola 68020, relied on external FPUs, such as the 68881 [14], similar to the way in which Intel’s pre-x486 and AMD’s pre-K5 architectures, used external units like the 8087. External FPUs were also seen in early SPARC systems and MIPS processors, such as the SPARCstation 1 with the Weitek 3170 FPU [15] and the MIPS R4000.
However, using co-processors introduced additional system complexity, requiring specialized hardware and coordination between the CPU and the floating-point unit (FPU). Despite this, co-processors became a staple in high-performance systems, enabling faster scientific computing and simulations.
@@ III-D The GPU Revolution
GPUs were initially developed in the 1980s and 1990s to meet the growing demand for 2D and 3D graphics rendering. Early GPUs, produced by companies such as SGI (Silicon Graphics), 3dfx, NVIDIA, and ATI (later acquired by AMD), were primarily focused on enhancing the real-time rendering of images, textures, and geometry, especially in gaming and graphical user interfaces (GUIs). A significant milestone came in 1999 with the release of NVIDIA’s GeForce 256 [19], marketed as the first “GPU” capable of processing graphics independently from the CPU. This allowed the CPU to focus on other tasks while the GPU specialized in real-time rendering. GPUs handled critical graphics tasks such as vertex transformations, lighting calculations, and texture mapping, all essential for 3D graphics in gaming and multimedia.
The early 2000s saw another major shift with the introduction of programmable graphics processing units (GPUs). NVIDIA’s GeForce3, launched in 2001, was a key milestone in this revolution, offering programmable shaders that allowed developers to directly program the GPU for custom operations, including floating-point computations [20].
The mid-2000s marked a turning point for GPUs, as their parallel processing capabilities were broadly recognized as useful in scientific and engineering computing. Unlike CPUs, which are optimized for serial processing and excel at handling a few threads quickly, GPUs have hundreds or even thousands of cores optimized for parallel workloads, making them ideal for tasks that can be divided into many smaller operations that thousands of threads can execute. This made GPUs especially useful in HPC, where they could accelerate simulations and large-scale mathematical computations, as well as in machine learning and AI, where they became indispensable for training deep learning models.
@@
NVIDIA’s CUDA [21] (Compute Unified Device Architecture) platform, first introduced in 2006, enabled a broad developer base to harness GPU power for non-graphics workloads. As a result, GPUs, initially designed for rendering graphics, soon became indispensable for general-purpose computing tasks, particularly in scientific computing [22]. Breakthroughs in AI algorithms in the same timeframe [23, 24] rapidly led to leveraging GPUs outside of scientific computing. The GPU’s ability to perform parallel floating-point operations at high speeds, complemented by equally impressive bandwidth capabilities (see Table I), firmly established them as ideal for diverse tasks requiring massive computational power, from training deep learning models to running large-scale simulations in physics and biology.
Today, GPUs are widely used in fields as varied as engineering and scientific simulations, medical research, cryptocurrency mining, big data processing, finance, and countless AI applications, cementing their role as a critical tool for a wide range of computational tasks beyond graphics rendering.
@@ IV-A The Rise of AI and Reduced Precision
Artificial intelligence, particularly deep learning, has profoundly impacted floating-point computation. AI workloads are dominated by matrix multiplications and tensor operations which can tolerate lower precision without a significant loss in accuracy. As a result, reduced-precision floating-point types, such as FP16 (half-precision), BF16 (brain floating-point, also referred to as bfloat16), and, most recently, FP8 [3] have become increasingly popular in AI applications (see Figure 1).
@@
It should be noted that the move toward reduced precision capabilities poses challenges for applications that require higher accuracy, such as scientific simulations and financial modeling. While AI applications can often tolerate lower precision, many scientific fields depend on high-precision calculations to ensure the validity of their results.
@@ IV-B Mixed-Precision Computing
In our discussion of mixed-precision computation, we will explore an approach that involves using different levels of precision at various stages of the computation process for solving a dense system of equations. Specifically, one phase of the computation, where the highest level of accuracy is not critical, may be performed using a lower (or even fixed) precision, such as half-precision (FP16), to optimize performance and reduce resource consumption [28].
As the computation progresses, we will switch to a higher precision, such as single-precision (FP32) or double-precision (FP64), for parts of the process where increased accuracy is essential. This transition is necessary in phases where errors accumulated during earlier steps must be corrected or refined, ensuring that the final result meets the desired accuracy thresholds. The dynamic adjustment between lower and higher precision enables a balance between computational speed, energy efficiency, and numerical precision. By strategically employing mixed precision, we can achieve significant performance gains without compromising the overall accuracy of the computation.
Mixed-precision computing has proven especially useful in AI [29] and HPC [30]. For example, in deep learning, models can be trained using a combination of FP8, BF16, FP16, and FP32 (single precision) calculations, reducing the computational load without sacrificing accuracy. The same domain affords opportunities for finer-grained application of this technique. For example, forward propagation through a neural network typically requires higher precision for weights and activations. In contrast, gradients in the backward propagation (used for updating weights) require a higher dynamic range. This has led to the introduction of two different FP8 variants, E4M3 and E5M2 [3]. Mixed-precision algorithms must manage these precision transitions efficiently to maximize performance without sacrificing the quality of results.
@@
This technique has also been applied in scientific computing, in cases where certain phases of simulations can tolerate lower precision, allowing for faster computations (see Table II and Figure 4). Using error-correction techniques, such as stochastic rounding [31] and iterative refinement to prevent the propagation of errors, allows mixed-precision algorithms to maintain high accuracy even when using reduced precision for certain operations, making mixed-precision computing suitable for various scientific and AI applications.
Figure 4: Representative power consumption curves measured on an NVIDIA Ampere A100 GPU during the execution of two different equations solvers is shown. The blue line shows the FP64 LU solver (corresponds to ZGETRF & ZGETRS in LAPACK) while the green line shows the Tensor Core accelerated mixed-precision iterative refinement solver available in the cuSOLVER library, which relies on cuBLAS for Level 3 BLAS operations, both for a matrix size of 32000. The mixed-precision solver is 4.4 times faster and 5.8 times more power-efficient. Comparison with other GPU architectures can be found in Table II.
By optimizing precision based on task requirements, mixed-precision computing improves both performance and energy efficiency, making it an essential technique for modern floating-point computation.
@@ IV-C Emulation in Modern Systems
Emulation, once a necessity in the absence of dedicated floating-point hardware, has come back in modern systems. Emulation techniques have evolved in terms of their underlying methods, implementations, and the flexibility with which they leverage (plentiful) hardware resources [32, 33, 34]. This adaptability is particularly important in domains such as AI and HPC, where computational needs often exceed the capabilities of the more established, higher-precision hardware resources.
TABLE III: Preliminary HPL Performance and Efficiency Measurements on Blackwell B200 GPU comparing emulation with s=7 [33] to native FP64 [data subject to change]
@@ V-A Balancing Precision and Efficiency
This trade-off is particularly pronounced when considering fields like AI, where reduced precision is sufficient for many tasks, and scientific computing, where precision cannot be sacrificed. Today, the same systems often run applications of both stripes and the integration of AI methods and scientific computing is a burgeoning field of study [35]. As a result, system designers must carefully consider the needs of each application when selecting the appropriate precision format.
@@
Another challenge in floating-point computation is the tension between specialized hardware (such as custom FPUs with the ability to execute compound instructions) and general-purpose processors (CPUs). On the one hand, specialized hardware is optimized for specific tasks, such as matrix multiplications in AI, but may lack the flexibility needed for broader applications. General-purpose processors, on the other hand, can handle a wide range of tasks but may not be as efficient for specialized computations.
@@ VI-B Energy Efficiency in Floating-Point Hardware
Energy efficiency has become a critical consideration in the design of floating-point hardware, particularly in large-scale computing environments like data centers and supercomputers. As computational workloads continue to grow, the energy required to power these systems has become a significant constraint, approaching 40 MegaWatts. The centrality of this is reflected in the growing attention given to the Green500 list, described in Section II.
Introduced by NVIDIA with the Volta [45] architecture in 2017, Tensor Cores are specialized hardware designed to accelerate matrix multiply-accumulate (MMA) operations, which are critical for deep learning tasks such as matrix multiplication and convolution in neural networks. These cores are particularly efficient due to the use of complex instruction types optimized for high-throughput operations (Table I). Tensor Cores are available across many different precisions, but are highly utilized in AI workloads that make extensive use of dense matrix operations. This complex operation approach significantly speeds up computations and reduces energy requirements without compromising requisite accuracy (see Figure 4 and Table II).
Tensor Cores [46] excel at accelerating matrix multiplications ($A[m\times k]\cdot B[k\times n]$), a fundamental operation in neural networks. They process these operations in blocks (e.g., (m, n, k) = (16, 16, 8)) [47], which boosts throughput for critical tasks like forward and backpropagation. This allows neural network operations that typically require hundreds of regular GPU instructions to be executed with relatively few tensor operations executed in a fraction of the time. Furthermore, Tensor Cores enable massive parallelism by executing multiple floating-point operations simultaneously and in a systolic manner, resulting in highly efficient throughput for deep learning models such as convolutional neural networks (CNNs) [24] or transformers [48]. Additionally, their power efficiency is notable, as the ability of a single Tensor Core to perform operations that would require multiple steps on a traditional GPU reduces energy consumption—a key consideration for both chip-level performance limits and the capabilities of large-scale data centers.
Energy-efficient hardware is especially important in AI and HPC environments, where systems must process massive amounts of data while minimizing their environmental impact. The development of low-power FPUs and specialized processors has enabled these systems to meet the growing demand for computational power without exceeding energy constraints.
@@ VIII Hardware’s Impact on Software
The interaction between hardware and software is deeply interdependent, with hardware setting the constraints within which software must operate. As hardware evolves, it directly influences how software is designed, written, and optimized. This co-evolution drives both fields forward, as hardware improvements open new opportunities for software innovation while software demands push hardware development to new heights.
@@ VIII-B Instruction Sets and Architectures
On CPUs, the hardware architecture, such as x86, ARM, or RISC-V, determines the instruction sets (ISAs) available to software. Software must be compatible with the hardware’s architecture, which affects how low-level operations are performed and how efficiently the software can execute. The introduction of SIMD (Single Instruction, Multiple Data) and specialized floating-point instructions has allowed scientific and AI software to accelerate matrix operations and other floating-point-intensive tasks.
Similarly, on GPUs, the Parallel Thread Execution (PTX) [47] ISA is available to enable it as a computing device. PTX is a part of CUDA, the parallel computing platform developed by NVIDIA that enables GPUs to perform general-purpose computing tasks beyond graphics rendering. It allows developers to leverage the massive parallel processing power of GPUs by offloading compute-intensive tasks from the CPU. CUDA provides a unified architecture that manages thousands of threads in a SIMT model (Single Instruction, Multiple Threads), supports various precision modes (such as FP16 and FP64), and optimizes memory access, making it highly efficient for parallel workloads like scientific computing, AI, and machine learning.
As NVIDIA’s GPUs have evolved, CUDA has remained central to utilizing them to maximum advantage, particularly with the introduction of Tensor Cores for mixed-precision operations, beginning with the Volta architecture. These innovations have led to dramatic increases in compute throughput, especially in AI/ML applications (Table I). Memory bandwidth has also significantly improved, offering higher Bytes/FLOP for non-Tensor Core operations. CUDA’s ability to manage parallel execution, memory hierarchy, and precision makes it essential for extracting maximum performance from modern GPUs.
@@ XI-A Impact of Algorithmic Complexity on Floating-Point Performance
Figure 5: Comparison of Bytes/FLOP across four generations of GPUs for FMA and Tensor Core throughput from Table I. The dashed line shows Tensor Core accelerated DGEMM performance using integer based emulation with 7 slices.

### v2

@@ II Benchmarks
Throughout this paper, we will refer to several community benchmarks that have emerged over time, each serving a critical role in evaluating and exposing performance characteristics and limitations of the underlying hardware, as well as providing a readily conveyed, widely understood record of progress. These benchmarks have become standard tools in the high-performance computing (HPC) community for assessing the efficiency and effectiveness of various computing systems. While in this paper, we focus on floating-point operation-focused benchmarks, other benchmarks exist. One such example is the Graph 500 [4], which measures a system’s performance on graph-based problems important for large dataset analysis.
@@
A more recent addition, HPL-MxP (formerly called HPL-AI) [10], or HPL for mixed-precision, is designed to benchmark systems optimized for AI and machine learning workloads. By focusing on mixed-precision operations, HPL-MxP reflects the growing need for systems capable of efficiently handling lower-precision calculations typical in AI models, exposing the performance capabilities of modern hardware in these emerging domains without abandoning the needs of scientific computing.
@@ III-B The Co-Processor Era
While the 8087 might be the best-known example of this technology, co-processors were in widespread use at the time. For example, the Motorola 68020, relied on external floating-point units (FPUs), such as the 68881 [14], similar to the way in which Intel’s pre-x486 and AMD’s pre-K5 architectures, used external units like the 8087. External FPUs were also seen in early SPARC systems and MIPS processors, such as the SPARCstation 1 with the Weitek 3170 FPU [15] and the MIPS R4000’s CP1 [16].
However, using co-processors introduced additional system complexity, requiring specialized hardware and coordination between the CPU and the FPU. Despite this, co-processors became a staple in high-performance systems, enabling faster scientific computing and simulations.
@@ III-D The GPU Revolution
GPUs (Graphics Processing Units) were initially developed in the 1980s and 1990s to meet the growing demand for 2D and 3D graphics rendering. Early GPUs, produced by companies such as SGI (Silicon Graphics), 3dfx, NVIDIA, and ATI (later acquired by AMD), were primarily focused on enhancing the real-time rendering of images, textures, and geometry, especially in gaming and graphical user interfaces (GUIs). A significant milestone came in 1999 with the release of NVIDIA’s GeForce 256 [20], marketed as the first “GPU” capable of processing graphics independently from the CPU. This allowed the CPU to focus on other tasks while the GPU specialized in real-time rendering. GPUs handled critical graphics tasks such as vertex transformations, lighting calculations, and texture mapping, all essential for 3D graphics in gaming and multimedia.
The early 2000s saw another major shift with the introduction of programmable GPUs. NVIDIA’s GeForce3, launched in 2001, was a key milestone in this revolution, offering programmable shaders that allowed developers to directly program the GPU for custom operations, including floating-point computations [21].
The mid-2000s marked a turning point for GPUs, as their parallel processing capabilities were broadly recognized as useful in scientific and engineering computing. Unlike CPUs, which are optimized for serial processing and excel at handling a few threads quickly, GPUs have hundreds or even thousands of cores optimized for parallel workloads, making them ideal for tasks that can be divided into many smaller operations that thousands of threads can execute. This made GPUs especially useful for HPC, where they could accelerate simulations and large-scale mathematical computations, and, later, for machine learning and AI, where they became indispensable for training deep learning models.
@@
NVIDIA’s CUDA [22] (Compute Unified Device Architecture) platform, first introduced in 2006, enabled a broad developer base to harness GPU power for non-graphics workloads. As a result, GPUs, initially designed for rendering graphics, soon became indispensable for general-purpose computing tasks, particularly in scientific computing [23, 24, 25]. Breakthroughs in AI algorithms came soon after [26, 27] and rapidly led to leveraging GPUs outside of scientific computing. The GPU’s ability to perform parallel floating-point operations at high speeds, complemented by equally impressive bandwidth capabilities (see Table I), firmly established them as ideal for diverse tasks requiring massive computational power, from training deep learning models to running large-scale simulations in physics and biology.
Today, GPUs are widely used in fields as varied as engineering and scientific simulations, medical research, cryptocurrency mining, big data processing, finance, and countless AI applications, cementing their role as a critical tool for a wide range of performance-critical computational tasks beyond graphics rendering.
@@ IV-A The Rise of AI and Reduced Precision
Artificial intelligence, particularly deep learning, has profoundly impacted floating-point computation. AI workloads are dominated by matrix multiplications and tensor operations which can tolerate lower precision without a significant loss in accuracy. As a result, reduced-precision floating-point types, such as FP16 (half-precision), BF16 (brain floating-point, also referred to as bfloat16), and, more recently, FP8 [3, 28, 29] have become increasingly popular in AI applications (see Figure 1).
@@
It should be noted that the move toward reduced-precision capabilities poses challenges for applications that require high accuracy, such as scientific simulations and financial modeling. While AI applications can often tolerate lower precision, many scientific fields depend on high-precision calculations to ensure the validity of their results.
@@ IV-B Mixed-Precision Computing
In the discussion of mixed-precision computation that follows, an approach is explored that involves using different levels of precision at various stages of the computation process for solving a dense system of equations. Specifically, one phase of the computation, where the highest level of accuracy is not critical, may be performed using a lower (or even fixed) precision, such as half-precision (FP16), to optimize performance and reduce resource consumption [33].
As the computation progresses, the algorithm switches to a higher precision, such as single-precision (FP32) or double-precision (FP64), to complete parts of the process where increased accuracy is essential. This transition is necessary in phases where errors accumulated during earlier steps must be corrected or refined, ensuring that the final result meets the desired accuracy thresholds. The dynamic adjustment between lower and higher precision enables a balance between computational speed, energy efficiency, and numerical precision. By strategically employing mixed precision, we can achieve significant performance gains without compromising the overall accuracy of the computation.
Mixed-precision computing has proven especially useful in AI [34] and HPC [35]. For example, in deep learning, models can be trained using a combination of FP8, BF16, FP16, and FP32 calculations, reducing the computational load without sacrificing accuracy. The same domain affords opportunities for finer-grained application of this technique. For example, forward propagation through a neural network typically requires higher precision for weights and activations. In contrast, gradients in the backward propagation (used for updating weights) require a higher dynamic range. This has led to the introduction of two different FP8 variants, E4M3 and E5M2 [3]. Mixed-precision algorithms must manage these precision transitions efficiently to maximize performance without forfeiting the quality of results.
@@
This technique has also been applied in scientific computing, in cases where certain phases of simulations can tolerate lower precision, allowing for faster computations (see Table II and Figure 4). Using error-correction techniques, such as stochastic rounding [36] and iterative refinement to prevent the propagation of errors, allows mixed-precision algorithms to maintain high accuracy even when using reduced precision for certain operations, making mixed-precision computing suitable for improving the performance of various scientific and AI applications.
Figure 4: Representative power consumption curves measured on an NVIDIA Ampere A100 GPU during the execution of two different equations solvers is shown. The blue line shows the FP64 LU solver (corresponds to ZGETRF & ZGETRS in LAPACK) while the green line shows the Tensor Core accelerated mixed-precision iterative refinement solver available in the cuSOLVER library, which relies on cuBLAS for Level 3 BLAS operations, both for a matrix size of 32000. Comparison with other GPU architectures can be found in Table II.
By optimizing precision based on task requirements, mixed-precision computing also improves energy efficiency, which is a critical constraint for modern floating-point computation. Figure 4 shows an example of a mixed-precision solver that is 4.4 times faster and 5.8 times more power-efficient than one using full precision for the entire algorithm.
@@ IV-C Emulation in Modern Systems
Emulation, once a necessity in the absence of dedicated floating-point hardware, has come back in modern systems. Emulation techniques have evolved in terms of their underlying methods, implementations, and the flexibility with which they leverage (plentiful) hardware resources [37, 38, 39]. This adaptability is particularly important in domains such as scientific AI and HPC, where computational needs often exceed the capabilities of the more established, higher-precision hardware resources.
TABLE III: Preliminary HPL Performance and Efficiency Measurements on Blackwell B200 GPU comparing emulation with s=7 (INT8 data elements) [38] to native FP64 [data subject to change]
@@ V-A Balancing Precision and Efficiency
This trade-off is particularly pronounced when considering fields like AI, where reduced precision is sufficient for many tasks, and scientific computing, where the accuracy of results cannot be sacrificed. Today, the same systems often run applications of both stripes and the integration of AI methods and scientific computing is a burgeoning field of study [40]. As a result, system designers must carefully consider the needs of each application when selecting the appropriate precision format.
@@
Another challenge in floating-point computation is the tension between specialized hardware (such as custom FPUs with the ability to execute compound instructions) and general-purpose processors (CPUs). On the one hand, specialized hardware is optimized for specific tasks, such as matrix multiplications in AI, but may lack the flexibility needed for broader applications. General-purpose processors, on the other hand, can handle a wide range of tasks but are not as efficient for specialized computations.
@@ VI-B Energy Efficiency in Floating-Point Hardware
Energy efficiency has become a critical consideration in the design of floating-point hardware across a wide variety of computing environments, from large-scale data centers and supercomputers to battery-reliant mobile devices such as laptops and phones. As computational workloads continue to grow, the energy required to power these platforms has become a significant constraint, with some public supercomputers approaching 40 megawatts. The centrality of energy consumption is reflected in the growing attention given to the Green500 list, described in Section II.
Introduced by NVIDIA with the Volta [50] architecture in 2017, Tensor Cores are specialized hardware designed to accelerate matrix multiply-accumulate (MMA) operations, which are critical for deep learning tasks such as matrix multiplication and convolution in neural networks. These cores are particularly efficient due to the use of complex instruction types optimized for high-throughput operations (Table I). Tensor Cores are available across many different precisions and are highly utilized in AI workloads that make extensive use of dense matrix operations. This complex operation approach significantly speeds up computations and reduces energy requirements without compromising requisite accuracy (see Figure 4 and Table II).
Tensor Cores [51] excel at accelerating matrix multiplications ($A[m\times k]\cdot B[k\times n]$), a fundamental operation in neural networks. They process these operations in blocks (e.g., (m, n, k) = (16, 16, 8)) [52], which boosts throughput for critical tasks like forward and backpropagation. This allows neural network operations that typically require hundreds of regular GPU instructions to be executed with relatively few tensor operations executed in a fraction of the time. Furthermore, Tensor Cores enable massive parallelism by executing multiple floating-point operations simultaneously and in a systolic manner, resulting in high throughput for deep learning models such as convolutional neural networks (CNNs) [27] or transformers [53]. Additionally, their power efficiency is notable, as the ability of a single Tensor Core to perform a chain of operations that would require multiple traditional GPU instructions reduces energy consumption—a key consideration for both chip-level performance limits and the capabilities of large-scale data centers. Thus, the development of low-power FPUs and specialized processors has enabled systems to meet the growing demand for computational power without exceeding energy constraints.
@@ VIII Hardware’s Impact on Software
The interaction between hardware and software is deeply interdependent, with hardware setting the constraints within which software must operate. As hardware evolves, it directly influences how software is designed, written, and optimized. Co-evolution drives both fields forward, as hardware improvements open new opportunities for software innovation while software demands push hardware development to new heights.
@@ VIII-B Instruction Sets and Architectures
On CPUs, the hardware architecture, such as x86, ARM, or RISC-V, determines the instruction sets (ISAs) available to software. Software must be compatible with the hardware’s architecture, which affects how low-level operations are performed and how efficiently the software can execute. As an example, the introduction of SIMD (Single Instruction, Multiple Data) and specialized floating-point instructions has allowed scientific and AI software to accelerate matrix operations and other floating-point-intensive tasks.
Similarly, on GPUs, the Parallel Thread Execution (PTX) [52] ISA is the interface to enable it as a computing device. PTX is a part of CUDA, the parallel computing platform developed by NVIDIA that enables GPUs to perform general-purpose computing tasks beyond graphics rendering. It allows developers to leverage the massive parallel processing power of GPUs by offloading compute-intensive tasks from the CPU. CUDA provides a unified architecture that manages thousands of threads in a SIMT model (Single Instruction, Multiple Threads), supports various precision modes (such as FP16 and FP64), and optimizes memory access, making it highly efficient for parallel workloads like scientific computing, AI, and machine learning.
As NVIDIA’s GPUs have evolved, CUDA has remained central to utilizing them to maximum advantage, particularly with the introduction of Tensor Core instructions for mixed-precision operations, beginning with the Volta architecture. These innovations have led to dramatic increases in compute throughput, especially in AI/ML applications (Table I). Memory bandwidth has also significantly improved, offering higher Bytes/FLOP for non-Tensor Core operations. CUDA’s ability to manage parallel execution, memory hierarchy, and precision makes it essential for extracting maximum performance from modern GPUs.
@@ XI-A Impact of Algorithmic Complexity on Floating-Point Performance
Figure 5: Comparison of Bytes/FLOP across four generations of GPUs for FMA and Tensor Core throughput from Table I. The dashed line shows Tensor Core accelerated DGEMM performance using integer based emulation with 7 slices (INT8 data storage elements) [38].